# LaB-RAG: Label Boosted Retrieval Augmented Generation for Radiology Report Generation

### v1

@@ 5 AP AXIAL 10 LAO 15 (Empty String)
RGRG generates the “Findings” section of a report by combining individual sentences that describe specific anatomical regions. This is accomplished by training a model over a specialized dataset, derived from MIMIC-CXR, which pairs anatomical region bounding boxes with sentences from the corresponding report detailing those regions. Thus RGRG learns to extract localized image latent embeddings to generate sentences grounded in the specific anatomical feature. As RGRG’s language model is a decoder-only transformer, the image embeddings are prepended input as tokens. Unlike CheXagent, where individual anatomical regions must be prompted by the user, RGRG automatically selects relevant regions using a custom trained object detector model. As RGRG only generates the “Findings”, we evaluate using our test split subset for studies with an extractable “Findings” section.
S7 Supplemental Results
Experiment Section Variable bleu4 rougeL bertscore f1radgraph f1chexbert
Baselines Findings LaB-RAG 0.042 0.204 0.861 0.183 0.466
@@
CheXagent 0.060 0.209 0.864 0.176 0.397
CXRMate 0.053 0.212 0.862 0.191 0.422
Both LaB-RAG 0.042 0.204 0.861 0.183 0.466
CheXagent 0.093 0.263 0.870 0.272 0.415
CXRMate 0.087 0.253 0.866 0.271 0.438
Filter & Format Findings Standard RAG 0.043 0.201 0.858 0.183 0.409
Label Filter only 0.041 0.198 0.857 0.180 0.443
Label Format only 0.045 0.208 0.863 0.193 0.466
LaB-RAG 0.042 0.204 0.861 0.183 0.466
Impression Standard RAG 0.020 0.149 0.851 0.137 0.382
Label Filter only 0.023 0.152 0.851 0.144 0.438
Label Format only 0.033 0.172 0.858 0.152 0.446
LaB-RAG 0.032 0.168 0.856 0.146 0.450
Filter Findings No-filter 0.045 0.208 0.863 0.193 0.466
Exact 0.042 0.204 0.861 0.183 0.466
Partial 0.043 0.205 0.862 0.185 0.467
Impression No-filter 0.033 0.172 0.858 0.152 0.446
Exact 0.032 0.168 0.856 0.146 0.450
Partial 0.033 0.174 0.858 0.154 0.451
Format Findings Naive 0.041 0.198 0.857 0.180 0.443
Simple 0.042 0.204 0.861 0.183 0.466
Verbose 0.040 0.196 0.856 0.180 0.438
Instruct 0.039 0.195 0.855 0.180 0.433
Impression Naive 0.023 0.152 0.851 0.144 0.438
Simple 0.032 0.168 0.856 0.146 0.450
Verbose 0.013 0.147 0.847 0.134 0.400
Instruct 0.013 0.146 0.846 0.131 0.393
LLM Findings Mistral-v1 0.037 0.178 0.850 0.181 0.454
BioMistral 0.038 0.197 0.857 0.168 0.445
Mistral-v3 0.042 0.204 0.861 0.183 0.466
Impression Mistral-v1 0.017 0.124 0.838 0.139 0.446
BioMistral 0.029 0.160 0.851 0.131 0.402
Mistral-v3 0.032 0.168 0.856 0.146 0.450
Label Findings True 0.055 0.221 0.866 0.203 0.568
Predicted 0.042 0.204 0.861 0.183 0.466
Impression True 0.047 0.229 0.871 0.221 0.800
Predicted 0.032 0.168 0.856 0.146 0.450
Section Intersection Findings-Intersect 0.050 0.214 0.863 0.197 0.431
Impression-Intersect 0.039 0.182 0.860 0.151 0.430
Both 0.061 0.226 0.865 0.211 0.456
Table S10: Full experimental results. By default, LaB-RAG uses “Exact” filter, “Simple” format/prompt, Mistral-v3, and predicted labels.
Figure S3: Comparison of LaB-RAG to literature baselines. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of comparisons vs LaB-RAG. By default, LaB-RAG uses “Exact” filter, “Simple” format/prompt, Mistral-v3, and predicted labels.
Figure S4: Comparison of label filter and format components of LaB-RAG. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. “Standard RAG” uses “No-filter” filter and “Naive” format/prompt. “Label Filter only” uses “Exact” filter and “Naive” format/prompt. “Label Format only” uses “No-filter” filter and “Simple” format/prompt. LaB-RAG otherwise defaults to “Exact” filter, “Simple” format/prompt, Mistral-v3, and predicted labels.
Figure S5: Comparison of LaB-RAG label filters. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. LaB-RAG otherwise defaults to “Simple” format/prompt, Mistral-v3, and predicted labels.
Figure S6: Comparison of LaB-RAG label formats. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. LaB-RAG otherwise defaults to “Exact” filter, Mistral-v3, and predicted labels.
Figure S7: Comparison of different LLM modules for LaB-RAG. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. LaB-RAG otherwise defaults to “Exact” label filter, “Simple” format/prompt, and predicted labels.
Figure S8: Comparison of label quality’s effect on LaB-RAG. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. LaB-RAG otherwise defaults to “Exact” format, “Simple” format/prompt, and Mistral-v3.
Figure S9: Comparison of generating different report sections and comparison of joint section generation against literature models. Significantly different pairs by paired t-test annotated if $p<0.05$, Bonferroni correction applied for number of all pairwise comparisons per-plot. LaB-RAG otherwise defaults to “Exact” format, “Simple” format/prompt, Mistral-v3, and predicted labels.
Figure S10: Histogram of image similarity rank for top 5 studies selected for each inference image, after “Exact” label filtering or “Partial” label reordering. While similar images may share some labels with the inference image, they may not match exactly.

### v2

@@ B.5 Literature Model Inference
RGRG: generates the “Findings” section of a report by combining individual sentences that describe specific anatomical regions. This is accomplished by training a model over a specialized dataset, derived from MIMIC-CXR, which pairs anatomical region bounding boxes with sentences from the corresponding report detailing those regions. Thus RGRG learns to extract localized image latent embeddings to generate sentences grounded in the specific anatomical feature. As RGRG’s language model is a decoder-only transformer, the image embeddings are prepended input as tokens. Unlike CheXagent, where individual anatomical regions must be prompted by the user, RGRG automatically selects relevant regions using a custom trained object detector model. As RGRG only generates the “Findings”, we evaluate using our test split subset for studies with an extractable “Findings” section.
Excluded models: As discussed in Section 2, RRG models can be categorized by the report section they generate. LaB-RAG is able to generate any target section given a corresponding retrieval corpus and thus comparison to other models only depend on code and data availability. FactMM-RAG (Sun et al., 2024) and CXR-RePaiR-Gen (Ranjit et al., 2023) are also retrieval based methods, however they do not share code for reproducibility. While BioViL-T (Bannur et al., 2023) and BioViL (Boecking et al., 2022) report RRG metrics and are usable as encoders, they do not share code for autoregressive or precise retrieval based generation. Finally, the critically missing component of Pragmatic Retrieval/Llama’s (Nguyen et al., 2023) codebase is tooling to extract the required “Indication” section. Nguyen et al. (2023) refer to the MIMIC Code Repository (Johnson et al., 2018) for this extraction, though we found the tool does not derive the “Indication” section. It is unclear to what precise modification is necessary to replicate the method of Nguyen et al. (2023).
Appendix C Label Format & Prompts
Table 4: The “Naive” label format does not incorporate labels.
You are an expert radiological assistant.
Your task is to generate a radiology report after <<Report>> given context information.
The context information contains examples of reports written for similar cases.
Use the examples to generate a report for the current case.
Strictly follow the instructions below to generate the reports.
*Instructions**
1. The report must be based on the information in the context.
2. The report must mimic the style of the reports shown in the context.
3. Do not generate blank reports.
CONTEXT:
Example: 1
(Report Text)
(More Examples)
Now generate the report for the current case.
Always generate reports based on the examples shown.
<<Report>>
Table 5: The “Simple” label format uses positive labels for the examples and target image.
You are an expert radiological assistant.
Your task is to generate a radiology report after <<Report>> given context information.
The context information contains examples of reports written for similar cases
and their associated labels.
Use the examples and their associated labels to generate a report for the current
case based on the current label.
Strictly follow the instructions below to generate the reports.
*Instructions**
1. The report must be based on the information in the context and the current label.
2. The report must mimic the style of the reports shown in the context.
3. Do not generate blank reports.
CONTEXT:
Example: 1
Label: (Positive Labels)
(Report Text)
(More Examples)
Now generate the report for the current case using its label below.
Always generate reports based on the examples shown.
Label: (Positive Labels)
<<Report>>
Table 6: The “Verbose” label format uses positive, negative, uncertain, and unmentioned labels for the examples and target image.
You are an expert radiological assistant.
Your task is to generate a radiology report after <<Report>> given context information.
The context information contains examples of reports written for similar cases
and their associated labels.
The labels provided are expert annotations.
More information about the labels is described below.
The individual labels used represent common chest radiographic observations
and fall under four categories: ‘Positive’, ‘Negative’, ‘Uncertain’ and ‘Unmentioned’.
These categories correspond to the mention or presence of the labels or their equivalent in the report.
Below is a description and example of each of the label categories:
1. ‘Positive’: A label is positive if the associated observation or disease is stated as present
       in the report, for example: ‘moderate bilateral effusions observed’.
2. ‘Negative’: A label is negative if the associated observation or disease is stated as absent
       in the report, for example: ‘no evidence of pulmonary edema’.
3. ‘Uncertain’: A label is uncertain if there is ambiguity about the presence or absence of
       the associated observation or disease in the report, for example: ‘pneumonia cannot be excluded
       in the appropriate clinical context’.
4. ‘Unmentioned’: A label is unmentioned if there is no mention of the associated observation
       or disease in report.
Use the examples, their associated labels, and the label descriptions to generate a report
for the current case based on the current label.
Strictly follow the instructions below to generate the reports.
*Instructions**
1. The report must be based on the information in the context and the current label.
2. The report must mimic the style of the reports shown in the context.
3. Do not generate blank reports.
CONTEXT:
Example: 1
Positive: (Positive Labels)
Negative: (Negative Labels)
Uncertain: (Uncertain Labels)
Unmentioned: (Unmentioned Labels)
(Report Text)
(More Examples)
Now generate the report for the current case using its label below.
Always generate reports based on the examples shown.
Positive: (Positive Labels)
Negative: (Negative Labels)
Uncertain: (Uncertain Labels)
Unmentioned: (Unmentioned Labels)
<<Report>>
Table 7: The “Instruct” label format uses the same format as “Verbose” with additional instructions.
(Same as Verbose)
*Instructions**
1. The report must be based on the information in the context and the current label.
2. The report must mimic the style of the reports shown in the context.
3. Do not generate blank reports.
4. Ensure that the positive labels are clearly described as being present in the report,
       using example language from the context.
5. Ensure that the negative labels are clearly described as being absent in the report,
       using example language from the context.
6. Describe the uncertain labels as necessary.
7. Ensure that the unmentioned labels are not mentioned in the report.
(Same as Verbose)
Appendix D Extended Tables
Table 8: Counts for selected single image view for each imaging study. Views presented in order of selection preference (e.g. PA before AP).
Dataset Section View, N (%) Overall Train Validate Test
MIMIC-CXR
Findings PA 71455 (45.9) 70204 (46.1) 546 (45.7) 705 (30.0)
AP 78015 (50.1) 75905 (49.9) 605 (50.6) 1505 (64.1)
LATERAL 160 (0.1) 156 (0.1) 2 (0.2) 2 (0.1)
$L$ 1396 (0.9) 1356 (0.9) 9 (0.8) 31 (1.3)
AP AXIAL 1 (0.0) 1 (0.0)
LAO 1 (0.0) 1 (0.0)
LPO 1 (0.0) 1 (0.0)
Unknown 4660 (3.0) 4522 (3.0) 34 (2.8) 104 (4.4)
Total 155689 152146 1196 2347
Impression PA 77755 (41.0) 76449 (41.1) 594 (39.1) 712 (32.0)
AP 104960 (55.4) 102697 (55.3) 877 (57.7) 1386 (62.3)
LATERAL 168 (0.1) 165 (0.1) 2 (0.1) 1 (0.0)
$L$ 1457 (0.8) 1420 (0.8) 8 (0.5) 29 (1.3)
AP AXIAL 1 (0.0) 1 (0.0)
LAO 1 (0.0) 1 (0.0)
Unknown 5210 (2.7) 5075 (2.7) 39 (2.6) 96 (4.3)
Total 189552 185808 1520 2224
CheXpert Plus
Findings PA 8568 (18.3) 8504 (18.3) 56 (17.7) 8 (13.1)
AP 38178 (81.6) 37865 (81.6) 260 (82.3) 53 (86.9)
Lateral 13 (0.0) 13 (0.0)
Total 46759 46382 316 61
Impression PA 28711 (15.3) 28495 (15.3) 185 (15.0) 31 (15.5)
AP 158823 (84.7) 157602 (84.7) 1052 (85.0) 169 (84.5)
Lateral 36 (0.0) 36 (0.0)
Total 187570 186133 1237 200
Table 9: Per-study demographics of patient samples.
Section Findings Impression
Split Overall Train Validate Test Overall Train Validate Test
MIMIC-CXR
Count, N Studies 155689 152146 1196 2347 189552 185808 1520 2224
Patients 60542 59794 459 289 62702 61935 479 288
Age Median [Q1, Q3] 64 [51,76] 64 [51,76] 62 [52,77] 69 [61,78] 65 [52,76] 64 [52,76] 62 [52,76] 69 [61,77]
Missing, N (%) 8132 (5.2) 7923 (5.2) 86 (7.2) 123 (5.2) 9829 (5.2) 9589 (5.2) 120 (7.9) 120 (5.4)
Sex, N (%) Female 73993 (47.5) 72425 (47.6) 538 (45.0) 1030 (43.9) 88578 (46.7) 86974 (46.8) 593 (39.0) 1011 (45.5)
Male 73564 (47.3) 71798 (47.2) 572 (47.8) 1194 (50.9) 91145 (48.1) 89245 (48.0) 807 (53.1) 1093 (49.1)
Unknown 8132 (5.2) 7923 (5.2) 86 (7.2) 123 (5.2) 9829 (5.2) 9589 (5.2) 120 (7.9) 120 (5.4)
Race or
Ethnicity, N (%)
White 89343 (57.4) 87143 (57.3) 729 (61.0) 1471 (62.7) 111465 (58.8) 109215 (58.8) 894 (58.8) 1356 (61.0)
Black 24720 (15.9) 24109 (15.8) 133 (11.1) 478 (20.4) 28726 (15.2) 28090 (15.1) 154 (10.1) 482 (21.7)
Hispanic/Latino 8641 (5.6) 8487 (5.6) 78 (6.5) 76 (3.2) 9980 (5.3) 9815 (5.3) 79 (5.2) 86 (3.9)
Asian 4810 (3.1) 4702 (3.1) 23 (1.9) 85 (3.6) 5835 (3.1) 5725 (3.1) 40 (2.6) 70 (3.1)
AIAN 351 (0.2) 343 (0.2) 4 (0.3) 4 (0.2) 506 (0.3) 492 (0.3) 5 (0.3) 9 (0.4)
NHPI 126 (0.1) 125 (0.1) 1 (0.1) 157 (0.1) 156 (0.1) 1 (0.1)
Other 4711 (3.0) 4570 (3.0) 56 (4.7) 85 (3.6) 5352 (2.8) 5145 (2.8) 118 (7.8) 89 (4.0)
Unknown 22987 (14.8) 22667 (14.9) 172 (14.4) 148 (6.3) 27531 (14.5) 27170 (14.6) 229 (15.1) 132 (5.9)
CheXpert Plus
Count, N Studies 46759 46382 316 61 187570 186133 1237 200
Patients 26695 26466 168 61 64702 64102 400 200
Age Median [Q1, Q3] 63 [50,75] 63 [50,75] 61 [51,73] 67 [55,77] 62 [49,74] 62 [49,74] 62 [51,73] 62 [48,74]
Missing, N (%) 59 (0.1) 59 (0.1) 214 (0.1) 212 (0.1) 2 (1.0)
Sex, N (%) Female 19233 (41.1) 19056 (41.1) 150 (47.5) 27 (44.3) 78063 (41.6) 77360 (41.6) 609 (49.2) 94 (47.0)
Male 27467 (58.7) 27267 (58.8) 166 (52.5) 34 (55.7) 109292 (58.3) 108560 (58.3) 628 (50.8) 104 (52.0)
Unknown 59 (0.1) 59 (0.1) 215 (0.1) 213 (0.1) 2 (1.0)
Race or
Ethnicity, N (%)
White 25709 (55.0) 25503 (55.0) 171 (54.1) 35 (57.4) 101939 (54.3) 101139 (54.3) 692 (55.9) 108 (54.0)
Black 2421 (5.2) 2404 (5.2) 15 (4.7) 2 (3.3) 9739 (5.2) 9655 (5.2) 76 (6.1) 8 (4.0)
Hispanic/Latino 6007 (12.8) 5967 (12.9) 31 (9.8) 9 (14.8) 23601 (12.6) 23454 (12.6) 122 (9.9) 25 (12.5)
Asian 5297 (11.3) 5255 (11.3) 34 (10.8) 8 (13.1) 19661 (10.5) 19529 (10.5) 108 (8.7) 24 (12.0)
AIAN 79 (0.2) 79 (0.2) 319 (0.2) 308 (0.2) 11 (0.9)
NHPI 663 (1.4) 654 (1.4) 7 (2.2) 2 (3.3) 2298 (1.2) 2274 (1.2) 19 (1.5) 5 (2.5)
Other 2023 (4.3) 2003 (4.3) 17 (5.4) 3 (4.9) 7561 (4.0) 7492 (4.0) 60 (4.9) 9 (4.5)
Unknown 4560 (9.8) 4517 (9.7) 41 (13.0) 2 (3.3) 22452 (12.0) 22282 (12.0) 149 (12.0) 21 (10.5)
Table 10: Per-study positive label prevalence.
Section Findings Impression
Split Overall Train Validate Test Overall Train Validate Test
Dataset Label, N (%)
MIMIC-CXR
Atelectasis 43504 (27.9) 42392 (27.9) 321 (26.8) 791 (33.7) 42728 (22.5) 41835 (22.5) 363 (23.9) 530 (23.8)
Cardiomegaly 43327 (27.8) 42002 (27.6) 319 (26.7) 1006 (42.9) 35041 (18.5) 34200 (18.4) 318 (20.9) 523 (23.5)
Consolidation 8885 (5.7) 8590 (5.6) 73 (6.1) 222 (9.5) 11716 (6.2) 11426 (6.1) 100 (6.6) 190 (8.5)
Edema 21606 (13.9) 20833 (13.7) 173 (14.5) 600 (25.6) 31991 (16.9) 31016 (16.7) 296 (19.5) 679 (30.5)
Enl. Card. 28858 (18.5) 27981 (18.4) 229 (19.1) 648 (27.6) 11236 (5.9) 10998 (5.9) 98 (6.4) 140 (6.3)
Fracture 6154 (4.0) 5998 (3.9) 29 (2.4) 127 (5.4) 3453 (1.8) 3388 (1.8) 17 (1.1) 48 (2.2)
Lung Lesion 6495 (4.2) 6271 (4.1) 67 (5.6) 157 (6.7) 6246 (3.3) 6080 (3.3) 59 (3.9) 107 (4.8)
Lung Opacity 41966 (27.0) 40771 (26.8) 304 (25.4) 891 (38.0) 36477 (19.2) 35626 (19.2) 284 (18.7) 567 (25.5)
No Finding 35201 (22.6) 34784 (22.9) 273 (22.8) 144 (6.1) 69600 (36.7) 68583 (36.9) 538 (35.4) 479 (21.5)
Pleural Effusion 37128 (23.8) 35918 (23.6) 301 (25.2) 909 (38.7) 45546 (24.0) 44443 (23.9) 400 (26.3) 703 (31.6)
Pleural Other 3903 (2.5) 3770 (2.5) 38 (3.2) 95 (4.0) 2247 (1.2) 2167 (1.2) 19 (1.2) 61 (2.7)
Pneumonia 14686 (9.4) 14262 (9.4) 114 (9.5) 310 (13.2) 26958 (14.2) 26284 (14.1) 207 (13.6) 467 (21.0)
Pneumothorax 5318 (3.4) 5208 (3.4) 32 (2.7) 78 (3.3) 7345 (3.9) 7245 (3.9) 50 (3.3) 50 (2.2)
Support Devices 44042 (28.3) 42772 (28.1) 322 (26.9) 948 (40.4) 44438 (23.4) 43581 (23.5) 410 (27.0) 447 (20.1)
Total 155689 152146 1196 2347 189552 185808 1520 2224
CheXpert Plus
Atelectasis 14581 (31.2) 14471 (31.2) 90 (28.5) 20 (32.8) 59736 (31.8) 59281 (31.8) 394 (31.9) 61 (30.5)
Cardiomegaly 10209 (21.8) 10126 (21.8) 67 (21.2) 16 (26.2) 29175 (15.6) 28957 (15.6) 194 (15.7) 24 (12.0)
Consolidation 8369 (17.9) 8295 (17.9) 63 (19.9) 11 (18.0) 35261 (18.8) 34978 (18.8) 253 (20.5) 30 (15.0)
Edema 13392 (28.6) 13270 (28.6) 108 (34.2) 14 (23.0) 60611 (32.3) 60131 (32.3) 426 (34.4) 54 (27.0)
Enl. Card. 8879 (19.0) 8807 (19.0) 62 (19.6) 10 (16.4) 18828 (10.0) 18675 (10.0) 143 (11.6) 10 (5.0)
Fracture 2664 (5.7) 2653 (5.7) 10 (3.2) 1 (1.6) 7395 (3.9) 7356 (4.0) 28 (2.3) 11 (5.5)
Lung Lesion 2809 (6.0) 2785 (6.0) 21 (6.6) 3 (4.9) 8131 (4.3) 8066 (4.3) 55 (4.4) 10 (5.0)
Lung Opacity 27126 (58.0) 26916 (58.0) 181 (57.3) 29 (47.5) 91069 (48.6) 90388 (48.6) 612 (49.5) 69 (34.5)
No Finding 1745 (3.7) 1731 (3.7) 11 (3.5) 3 (4.9) 15677 (8.4) 15532 (8.3) 110 (8.9) 35 (17.5)
Pleural Effusion 22025 (47.1) 21855 (47.1) 146 (46.2) 24 (39.3) 84656 (45.1) 84010 (45.1) 583 (47.1) 63 (31.5)
Pleural Other 1568 (3.4) 1557 (3.4) 8 (2.5) 3 (4.9) 4562 (2.4) 4534 (2.4) 25 (2.0) 3 (1.5)
Pneumonia 3742 (8.0) 3709 (8.0) 30 (9.5) 3 (4.9) 19917 (10.6) 19760 (10.6) 136 (11.0) 21 (10.5)
Pneumothorax 6207 (13.3) 6164 (13.3) 36 (11.4) 7 (11.5) 18097 (9.6) 17978 (9.7) 99 (8.0) 20 (10.0)
Support Devices 28157 (60.2) 27922 (60.2) 202 (63.9) 33 (54.1) 106210 (56.6) 105420 (56.6) 695 (56.2) 95 (47.5)
Total 46759 46382 316 61 187570 186133 1237 200
Table 11: Test-split per-label classifier F1 scores. By default, LaB-RAG uses classifiers trained on CheXbert extracted labels and dataset adapted embeddings.
Model Labeler Dataset Section Label
Atelectasis Cardiomegaly Consolidation Edema Enl. Card. Fracture Lung Lesion
biovilt chexbert mimic-cxr findings 0.59 0.71 0.10 0.62 0.46 0.09 0.04
impression 0.50 0.06 0.00 0.68 0.20 0.04 0.04
chexpert mimic-cxr findings 0.53 0.67 0.11 0.53 0.03 0.09 0.04
impression 0.42 0.07 0.01 0.59 0.05 0.03 0.15
gloria chexbert chexpertplus findings 0.00 0.11 0.00 0.61 0.30 0.00 0.00
impression 0.00 0.46 0.00 0.70 0.19 0.25 0.00
chexpert chexpertplus findings 0.00 0.11 0.00 0.62 0.00 0.00 0.00
impression 0.41 0.48 0.00 0.63 0.20 0.35 0.00
resnet50 chexbert chexpertplus findings 0.46 0.00 0.00 0.67 0.32 0.00 0.10
impression 0.48 0.08 0.35 0.60 0.00 0.29 0.00
mimic-cxr findings 0.55 0.66 0.19 0.57 0.44 0.02 0.16
impression 0.00 0.00 0.21 0.63 0.03 0.06 0.11
Lung Opacity No Finding Pleural Effusion Pleural Other Pneumonia Pneumothorax Support Devices
biovilt chexbert mimic-cxr findings 0.59 0.35 0.73 0.04 0.04 0.12 0.79
impression 0.45 0.57 0.70 0.08 0.01 0.07 0.64
chexpert mimic-cxr findings 0.67 0.37 0.70 0.00 0.04 0.08 0.82
impression 0.49 0.56 0.68 0.10 0.07 0.07 0.65
gloria chexbert chexpertplus findings 0.75 0.40 0.80 0.00 0.00 0.40 0.83
impression 0.60 0.50 0.77 0.00 0.34 0.47 0.87
chexpert chexpertplus findings 0.79 0.40 0.68 0.00 0.00 0.38 0.86
impression 0.58 0.50 0.00 0.00 0.00 0.50 0.83
resnet50 chexbert chexpertplus findings 0.71 0.00 0.60 0.80 0.00 0.00 0.74
impression 0.57 0.40 0.54 0.00 0.00 0.00 0.71
mimic-cxr findings 0.02 0.26 0.64 0.04 0.00 0.10 0.68
impression 0.00 0.42 0.59 0.05 0.02 0.14 0.56
Table 12: Full experiment results on MIMIC-CXR. BL-4: BLEU-4, RG-L: ROUGE-L, BERT: BERTScore, F1-RG: F1-RadGraph, F1-CXB: F1-CheXbert.
Section Findings Impression
Metric BL-4 RG-L BERT F1-RG F1-CXB BL-4 RG-L BERT F1-RG F1-CXB
Experiment Variable
Literature LaB-RAG 0.042 0.205 0.861 0.187 0.524 0.031 0.166 0.856 0.144 0.439
@@ Appendix D Extended Tables
Core Standard RAG 0.043 0.201 0.858 0.183 0.409 0.020 0.149 0.851 0.137 0.382
Label Filter only 0.042 0.202 0.858 0.188 0.504 0.022 0.152 0.851 0.138 0.425
Label Format only 0.043 0.207 0.861 0.190 0.497 0.030 0.170 0.857 0.148 0.432
LaB-RAG 0.042 0.205 0.861 0.187 0.524 0.031 0.166 0.856 0.143 0.438
Filter No-filter 0.043 0.207 0.861 0.190 0.497 0.030 0.170 0.857 0.148 0.432
Exact 0.042 0.205 0.861 0.187 0.524 0.031 0.166 0.856 0.143 0.438
Partial 0.042 0.205 0.861 0.189 0.513 0.029 0.166 0.857 0.145 0.445
Prompt Naive 0.042 0.202 0.858 0.188 0.504 0.022 0.152 0.851 0.138 0.425
Simple 0.042 0.205 0.861 0.187 0.524 0.031 0.166 0.856 0.143 0.438
Verbose 0.041 0.206 0.860 0.193 0.508 0.028 0.165 0.855 0.143 0.431
Instruct 0.041 0.204 0.859 0.192 0.501 0.025 0.162 0.852 0.144 0.427
Language Model Mistral-v1 0.035 0.173 0.846 0.188 0.516 0.016 0.123 0.838 0.133 0.430
BioMistral 0.038 0.199 0.859 0.172 0.509 0.028 0.162 0.852 0.130 0.414
Mistral-v3 0.042 0.205 0.861 0.187 0.524 0.031 0.166 0.856 0.143 0.438
Embedding Model BioViL-T 0.042 0.205 0.861 0.187 0.524 0.031 0.166 0.856 0.143 0.438
ResNet50 0.033 0.190 0.858 0.163 0.430 0.023 0.150 0.854 0.125 0.326
Label Quality Extracted - CheXbert 0.057 0.226 0.868 0.219 0.942 0.050 0.233 0.872 0.224 0.950
Extracted - CheXpert 0.056 0.223 0.867 0.211 0.741 0.047 0.229 0.871 0.221 0.801
Predicted - CheXbert 0.042 0.205 0.861 0.187 0.524 0.031 0.166 0.856 0.143 0.438
Predicted - CheXpert 0.042 0.205 0.862 0.186 0.496 0.032 0.172 0.857 0.150 0.438
Retrieved Samples 3 0.040 0.202 0.860 0.184 0.524 0.020 0.157 0.854 0.132 0.437
5 0.042 0.205 0.861 0.187 0.524 0.031 0.166 0.856 0.143 0.438
10 0.044 0.207 0.861 0.189 0.523 0.035 0.172 0.856 0.146 0.437
Table 13: Full experiment results on CheXpert Plus. BL-4: BLEU-4, RG-L: ROUGE-L, BERT: BERTScore, F1-RG: F1-RadGraph, F1-CXB: F1-CheXbert.
Section Findings Impression
Metric BL-4 RG-L BERT F1-RG F1-CXB BL-4 RG-L BERT F1-RG F1-CXB
Experiment Variable
Core Standard RAG 0.049 0.194 0.847 0.182 0.459 0.031 0.188 0.830 0.111 0.441
Label Filter only 0.038 0.202 0.848 0.204 0.441 0.030 0.197 0.831 0.116 0.509
Label Format only 0.040 0.191 0.847 0.192 0.478 0.046 0.213 0.839 0.159 0.502
LaB-RAG 0.033 0.202 0.849 0.200 0.451 0.043 0.215 0.837 0.149 0.505
Filter No-filter 0.040 0.191 0.847 0.192 0.478 0.046 0.213 0.839 0.159 0.502
Exact 0.033 0.202 0.849 0.200 0.451 0.043 0.215 0.837 0.149 0.505
Partial 0.034 0.196 0.847 0.195 0.471 0.041 0.212 0.838 0.152 0.503
Prompt Naive 0.038 0.202 0.848 0.204 0.441 0.030 0.197 0.831 0.116 0.509
Simple 0.033 0.202 0.849 0.200 0.451 0.043 0.215 0.837 0.149 0.505
Verbose 0.042 0.203 0.851 0.203 0.460 0.042 0.210 0.836 0.135 0.497
Instruct 0.041 0.207 0.850 0.194 0.436 0.040 0.202 0.834 0.130 0.488
Language Model Mistral-v1 0.026 0.150 0.828 0.191 0.426 0.030 0.175 0.826 0.134 0.492
BioMistral 0.035 0.193 0.847 0.180 0.417 0.010 0.154 0.816 0.076 0.372
Mistral-v3 0.033 0.202 0.849 0.200 0.451 0.043 0.215 0.837 0.149 0.505
Embedding Model GLoRIA 0.033 0.202 0.849 0.200 0.451 0.043 0.215 0.837 0.149 0.505
ResNet50 0.017 0.158 0.839 0.151 0.445 0.015 0.170 0.826 0.080 0.419
Label Quality Extracted - CheXbert 0.045 0.215 0.854 0.244 0.939 0.059 0.253 0.845 0.188 0.967
Extracted - CheXpert 0.046 0.214 0.852 0.225 0.721 0.059 0.248 0.842 0.187 0.815
Predicted - CheXbert 0.033 0.202 0.849 0.200 0.451 0.043 0.215 0.837 0.149 0.505
Predicted - CheXpert 0.045 0.209 0.850 0.203 0.457 0.035 0.199 0.833 0.121 0.488
Retrieved Samples 3 0.039 0.206 0.849 0.194 0.440 0.040 0.205 0.836 0.155 0.509
5 0.033 0.202 0.849 0.200 0.451 0.043 0.215 0.837 0.149 0.505
10 0.036 0.204 0.848 0.185 0.448 0.043 0.221 0.838 0.151 0.505
Appendix E Extended Figures
Figure 6: Image embedding similarity rank of label-filtered retrieved samples on MIMIC-CXR.
E.1 MIMIC-CXR Experiments
Figure 7: Comparison of LaB-RAG to literature models on MIMIC-CXR. CXR-RePaiR, CXR-ReDonE, and X-REM are other retrieval based models, like LaB-RAG. RGRG, CheXagent, and CXRMate are fine-tuned models.
Figure 8: Ablations of LaB-RAG’s core label filter and label format compared to standard RAG on MIMIC-CXR.
Figure 9: Variations of LaB-RAG label filter on MIMIC-CXR. By default, LaB-RAG uses the Exact filter.
Figure 10: Variations of LaB-RAG label format on MIMIC-CXR. By default, LaB-RAG uses the Simple format.
Figure 11: Alternate generative language models for LaB-RAG on MIMIC-CXR. By default, LaB-RAG uses Mistral-v3.
Figure 12: Alternate image embedding models for LaB-RAG on MIMIC-CXR. By default, LaB-RAG uses the dataset adapted model, in this case BioViL-T for MIMIC-CXR.
Figure 13: Ablation experiment testing the impact of label quality on LaB-RAG on MIMIC-CXR. Extracted labels are derived from the ground-truth report using either the CheXbert or CheXpert labelers. Predicted labels are inferred using linear classifiers trained over the respective label type. By default, LaB-RAG uses predicted CheXbert labels.
Figure 14: Variations on number of retrieved reports for LaB-RAG on MIMIC-CXR. By default, LaB-RAG uses 5 retrieved reports.
E.2 CheXpert Plus Experiments
Figure 15: Ablations of LaB-RAG’s core label filter and label format compared to standard RAG on CheXpert Plus.
Figure 16: Variations of LaB-RAG label filter on CheXpert Plus. By default, LaB-RAG uses the Exact filter.
Figure 17: Variations of LaB-RAG label format on CheXpert Plus. By default, LaB-RAG uses the Simple format.
Figure 18: Alternate generative language models for LaB-RAG on CheXpert Plus. By default, LaB-RAG uses Mistral-v3.
Figure 19: Alternate image embedding models for LaB-RAG on CheXpert Plus. By default, LaB-RAG uses the dataset adapted model, in this case GLoRIA for CheXpert Plus.
Figure 20: Ablation experiment testing the impact of label quality on LaB-RAG on CheXpert Plus. Extracted labels are derived from the ground-truth report using either the CheXbert or CheXpert labelers. Predicted labels are inferred using linear classifiers trained over the respective label type. By default, LaB-RAG uses predicted CheXbert labels.
Figure 21: Variations on number of retrieved reports for LaB-RAG on CheXpert Plus. By default, LaB-RAG uses 5 retrieved reports.